Teas costing $7 per pound are to be mixed with teas costing $4 per pound to make a 30-pound mixture. If the mixture is to sell for $6 per pound, how many pounds of each kind of tea should be used?
Answer by nerdybill(7384)

Teas costing $7 per pound are to be mixed with teas costing $4 per pound to make a 30-pound mixture. If the mixture is to sell for $6 per pound, how many pounds of each kind of tea should be used?
.
Let x = amount (lbs) of $7 tea
then
30-x = amount of $4 tea
.
7x + 4(30-x) = 6(30)
7x + 120-4x = 180
3x + 120 = 180
3x = 60
x = 20 lbs (of $7 tea)
.
amount of $4 tea:
30-x = 30-20 = 10 lbs